# Uniqueness of Galilean and Carrollian limits of gravitational theories and application to higher derivative gravity

### 01 Jan 2024 v1

@@ IV Applications to HDG
Having established that all methods of taking the Galilean and Carrollian limits are equivalent, we derive the $n-$th order of the Galilean and Carrollian expansions of a generic HDG theory.
@@ IV.1 $f(R)$ gravity
where $R_{1},R_{2},R_{3}$ are the expansion terms shown explicitly in eq.(3.2) in [58] and are displayed in (84). (The use of the connection does not affect the from of the expansion as shown in Sec. III, different connections will result in different forms for $R_{1}$,$R_{2}$,$R_{3}$.)
@@
where $a_{ij}$ are constants without any $c$ dependency. Expanding $f(R)$, we get
@@
In this case we need to be extra careful since $j$ can be unbounded regardless of the value of $i$, and appears in the power of $c$. For a theory to have consistent Galilean and Carrollian limits, the quantity ${2k-2n+j}$ must be bounded from above and below.
@@ IV.1.1 Carrollian limit
Let $l=\min(2k-2n+j)$. The electric limit’s Lagrangian is given by
@@
$=c^{l}\sum_{i=0}^{\infty}\sum_{j=-\infty}^{\infty}a_{ij}\sum_{\mathclap{\begin{subarray}{c}n,m\geq 0\\ 2n+m+\frac{l-j}{2}=i\end{subarray}}}\hskip 5.0pt\frac{i!}{n!m!\big{(}n+\frac{l-j}{2}\big{)}!}R_{1}^{n}R_{2}^{m}R_{3}^{n+\tfrac{l-j}{2}}.$
The magnetic limit’s Lagrangian is given by
@@
As an example we compute the Carrollian limits of the theories $f(R)=R+c^{N}\alpha R^{2}$ where $\alpha$ is a constant with no $c$ dependency. The Carrollian limit of these type of theories was computed in [77]. The electric limit can be computed from (43) by noticing that all the coefficients $a_{ij}$ are zero except $a_{10}=1$ and $a_{2N}=\alpha$. In this case $i=1,2$ and $j=0,N$.
@@
In order for the second term to couple to some higher order of the first term, we must set $N\geq 2$ otherwise, it will be pure quadratic gravity not a modification of GR. This is the same condition derived in [77]. The magnetic limit can be derived from (44) by noting that the first term has $i=1$, $j=0$, $l=-2$. Solving $2n+m+\frac{l+2-j}{2}=i$, we get $m=1$ and $n=0$. While the second term has $i=2$, $j=N$ and $l=N-4$. By solving $2n+m+\tfrac{l+2-j}{2}=i$, we get $m=n=1$. Putting all together we get
@@
As an example, we calculate the electric and magnetic limits of GR ($f(R)=R$). In this case $g_{i}=0$ except for $g_{1}=1$. Thus
$\mathcal{L}_{\textrm{el}}=c^{l}\sum_{\mathclap{\begin{subarray}{c}n,m\geq 0\\ 2n+m+\frac{l}{2}=1\end{subarray}}}\hskip 5.0pt\frac{1}{n!m!\big{(}\frac{l}{2}+n\big{)}!}R_{1}^{n}R_{2}^{m}R_{3}^{\tfrac{l}{2}+n},$ (49)
@@
The $n-$th order in the Carrollian expansion can be derived from the same minimization problem with replacing $l$ with $l+2n$.
@@ IV.1.2 Galilean limit
$=c^{l}\sum_{i=0}^{\infty}\sum_{j=-\infty}^{\infty}a_{ij}\sum_{\mathclap{\begin{subarray}{c}n,m\geq 0\\ 2n+m+\frac{l-j}{2}=i\end{subarray}}}\hskip 5.0pt\frac{i!}{n!m!\big{(}n+\frac{l-j}{2}\big{)}!}R_{1}^{n}R_{2}^{m}R_{3}^{n+\frac{l-j}{2}}.$
which is similar to the Carrollian case but with $l$ is a maximum rather than a minimum. The NLO is given by
@@
we will derive the Galilean limit of $f(R)=R+c^{N}\alpha R^{2}$ as an example. As mentioned in the previous section all coefficients $a_{ij}=0$ except $a_{10}=1$ and $a_{2N}=\alpha$. The first term in the Galilean limit has $i=1$, $j=0$ and $l=2$ i.e, $n=0$. Solving $2n+m+\tfrac{l-j}{2}=i$ gives $m=0$. The second term has $i=2$, $j=N$, $l=N+4$ i.e, $n=0$. Solving $2n+m+\tfrac{l-j}{2}=i$ gives $m=0$. Putting all together, we get
@@
For this theory to be a modification of the Galilean limit of GR at some order, we must impose the condition $N\leq-2$, otherwise it will be a theory of pure quadratic gravity.
@@
with the condition $N\leq-2$ to get a GR modification. From this condition and the condition for the Carrollian case, we see that no $f(R)$ theory with finite order can be a modification of GR in both the Galilean and the Carrollian regimes at the same time.
@@
The NLO, however, will be a little different, the general formula can be derived similar to the Carrollian case so we write the result directly
@@
We compute GR as an example: The first order can be derived from
@@
The $n-$th order in the Galilean expansion can be derived from the same maximization problem with replacing $l$ with $l-2n$.
@@ IV.2 $f(g_{\mu\nu},R_{\mu\nu\sigma\rho})$ gravity
We first consider gravity theories of the form $f(g_{\mu\nu},R_{\mu\nu\sigma\rho})$, in the paper’s notation the Lagrangian can be written as
@@
where $\mathcal{R}$ can be the Riemann tensor, the Ricci tensor with all indices down, or the Ricci scalar, $\bar{\mathcal{R}}$ can be the same tensors with all indices up, and $a,b$ are positive numbers such that the Lagrangian terms are scalars. Using the formulae in [77], we can write the generic expansions
$\mathcal{R}$ $=\tfrac{1}{c^{2}}\mathcal{R}_{1}+\mathcal{R}_{2}+c^{2}\mathcal{R}_{3}+c^{4}\mathcal{R}_{4},$$\hskip 10.0pt\bar{\mathcal{R}}$ $=\tfrac{1}{c^{4}}\bar{\mathcal{R}}_{1}+...+c^{2}\bar{\mathcal{R}}_{4}.$ (64)
The explicit expressions are in the appendix. Raising the expansions to the powers of $a$, $b$ and $c$ respectively, we get
@@
Assuming $\alpha=c^{N}\alpha^{\prime}$ where $\alpha^{\prime}$ has no $c$ dependency and substituting in the Lagrangian, we get
@@
The $n$-th order in the Galilean or the Carrollian limits can be obtained by solving a numerical optimization problem as we will show.
@@ IV.2.1 Galilean limit
This is a constrained numerical optimization problem that can be solved by the simplex method for $k_{i}$, $\bar{k}_{j}$ given that $a,b$ are known. The simplex method is explained in [80]. The solution is $z=4a+4b+2+N$. Thus, for this term to be a modification of GR, the condition on $N$ is $N\leq-4a-4b$. The $n$-th order can be deduced by modifying the constrained optimization problem by $z\to z-2n$.
@@ IV.2.2 Carrollian limit
$\sum_{i=0}^{4}k_{i}=a,$ $\hskip 10.0pt\sum_{j=0}^{4}\bar{k}_{j}=b,$ $k_{i},\bar{k}_{j},\geq 0,$ (70)
which is again solvable by the simplex method given that $a,b$ are known.The solution is $z=N-2a-4b-2$. Thus, for the term to be a modification of GR, the condition on $N$ must be $N\geq 2a+4b$. The $n$-th order is given by the modified optimization problem with $z\to z+2n$.
We can see that the conditions on $N$ in the two limits are mutually exclusive. Thus, no such theory (of the form (63)) with finite $a,b$ can be a GR modification in the two limits simultaneously. In both limits we can recover the results from the previous section if we set =¯$\mathcal{R}=\bar{\mathcal{R}}=R$. We can recover quadratic gravity examples in the previous section by setting $a=2,b=0$ or $a=0,b=2$ or $a=b=1$.
@@ IV.3 $f(g_{\mu\nu},R_{\mu\nu\sigma\rho},\nabla_{\mu})$ gravity
Now we consider the most general HDG theory of finite order. The most general Lagrangian was proposed in [81], and is given by
@@
where $\mathcal{P}_{0}$ and $\mathcal{Q}_{iI}$ are tensors made up of the Riemann tensor and the metric, and $O_{iI}$ are differential operators made up of contractions of Covariant derivatives. The first term is just $f(g_{\mu\nu},R_{\mu\nu\sigma\rho})$ which was discussed in the previous subsection. The second term of (71), in the notation we use in this paper, can be rearranged as
@@
where we added a constant with no $c$ dependency $\alpha^{\prime}$, and $i_{\textrm{max}}$, $j_{\textrm{max}}$, $a_{i}$, $b_{ij}$, $c_{ij}$ and $N$ are integers. Expanding the terms using (64) we get
@@
Putting all together, we get the action
@@
As the previous cases, the Galilean and Carrollian expansions are transformed into constrained optimization problems as follows:
@@ IV.3.2 Carrollian limit
We have shown that theories of the form (63) as well as quadratic gravity do not satisfy these conditions. We will leave the search for such theory (if any exists) to future work.
V Summary and conclusions
In this paper, we reviewed the methods used to construct non-Lorentzian gravitational theories from Lorentzian ones. We showed that all methods of taking the Galilean (or Carrollian) limit give the same metric, thus, the same non-Lorentzian theories. In the case of the Galilean expansion, the GAG procedure gives the same Galilean theories as the PNR approach. However, the PNR approach explore more theories from expanding the PNR quantities in powers of $\tfrac{1}{c^{2}}$. Such expansion gives rise to infinite number of theories, one for each order, which are not accounted for in other methods. However, these theories are not Galilean till the NLO. The same Galilean theories can be deduced by performing the ADM decomposition and taking the infinite limit of the norm of the normal vector. A similar situation occur in the Carrollain expansion, although the ZS approach (which give the same Carrollian theories as the Carroll group gauging approach) is computationally easier than the PUL approach, the PUL approach explore a larger space of theories, the Carrollian theories to the NLO coincide with the theories we get from other approaches.
Having established that all approaches give the same metric, we can write general expansions for the curvature tensors. We explored the most general finite order HDG theory introducing an algorithm to calculate the $n$-th order of its Galilean and Carrollian expansions as follows:
@@ V Summary and conclusions
    We write the Lagrangian and identify which form it takes (is it of the form (39), (63) or the general (74)) and if the Lagrangian depends on $c$ explicitly.
@@
Another interesting future direction of research is to search for gravity theories which are viable modifications of GR in both Carrollian, and Galilean limits at the same time i.e, satisfies the conditions (79). If such theory exists, it is interesting to see what significance it has to be such a modification. Is there a defining property of the theory that allows this? if so, how does it impact its solutions. It is also interesting to apply the algorithm introduced to study its Galilean and Carrollian limits, and see how does it modify black holes in GR.

### 23 Jan 2024 v2

@@ IV Applications to HDG
Having established that all methods of taking the Galilean and Carrollian limits are equivalent, we derive an algorithm to calculate the $n-$th order of the Galilean and Carrollian expansions of a generic HDG theory.
@@ IV.1 $f(R)$ gravity
where $R_{1},R_{2},R_{3}$ are the expansion terms shown explicitly in (3.2) in [64]; they are also displayed in (87). (The use of the connection does not affect the from of the expansion as shown in Sec. III, different connections will result in different forms for $R_{1}$,$R_{2}$,$R_{3}$.)
@@
where $a_{ij}$ are constants without any $c$ dependency, and expanding $f(R)$, we get
@@
Here, we need to be extra careful since $j$ can be unbounded regardless of the value of $i$, and appears in the power of $c$. For a theory to have consistent Galilean and Carrollian limits, the quantity ${2k-2n+j}$ must be bounded from above and below.
@@ IV.1.1 Carrollian limit
Let us denote $l=\min(2k-2n+j)$. The electric limit’s Lagrangian is given by
@@
$=c^{l}\sum_{i=0}^{\infty}\sum_{j=-\infty}^{\infty}a_{ij}\sum_{\mathclap{\begin{subarray}{c}n,m\geq 0\\ 2n+m+\frac{l-j}{2}=i\end{subarray}}}\hskip 5.0pt\frac{i!}{n!m!\big{(}n+\frac{l-j}{2}\big{)}!}R_{1}^{n}R_{2}^{m}R_{3}^{n+\tfrac{l-j}{2}},$
while the magnetic limit’s Lagrangian reads
@@
As a specific example we now compute the Carrollian limits of the theories $f(R)=R+c^{N}\alpha R^{2}$ where $\alpha$ is a constant with no $c$ dependency. The Carrollian limit of these type of theories was computed in [85]. The electric limit can be computed from (46) by noticing that all the coefficients $a_{ij}$ are zero except $a_{10}=1$ and $a_{2N}=\alpha$. In this case $i=1,2$ and $j=0,N$.
@@
In order for the second term to couple to some higher order of the first term, we must set $N\geq 2$ otherwise, it will be pure quadratic gravity, i.e., not a modification of GR. This is the same condition derived in [85]. The magnetic limit can be derived from (47) by noting that the first term has $i=1$, $j=0$, $l=-2$. Solving $2n+m+\frac{l+2-j}{2}=i$, we get $m=1$ and $n=0$. The second term has $i=2$, $j=N$ and $l=N-4$. Again, by solving $2n+m+\tfrac{l+2-j}{2}=i$, we get $m=n=1$. Putting all together we obtain
@@
As consistency check, let us calculate the electric and magnetic limits of GR, $f(R)=R$. In this case $g_{i}=0$ except for $g_{1}=1$; thus
$\mathcal{L}_{\textrm{el}}=c^{l}\sum_{\mathclap{\begin{subarray}{c}n,m\geq 0\\ 2n+m+\frac{l}{2}=1\end{subarray}}}\hskip 5.0pt\frac{1}{n!m!\big{(}\frac{l}{2}+n\big{)}!}R_{1}^{n}R_{2}^{m}R_{3}^{\tfrac{l}{2}+n}.$ (52)
@@
The $n-$th order in the Carrollian expansion can be derived from the same minimization problem by replacing $l$ with $l+2n$.
@@ IV.1.2 Galilean limit
$=c^{l}\sum_{i=0}^{\infty}\sum_{j=-\infty}^{\infty}a_{ij}\sum_{\mathclap{\begin{subarray}{c}n,m\geq 0\\ 2n+m+\frac{l-j}{2}=i\end{subarray}}}\hskip 5.0pt\frac{i!}{n!m!\big{(}n+\frac{l-j}{2}\big{)}!}R_{1}^{n}R_{2}^{m}R_{3}^{n+\frac{l-j}{2}},$
which is similar to the Carrollian case but with $l$ being the maximum rather than the minimum. The NLO reads
@@
We will derive the Galilean limit of $f(R)=R+c^{N}\alpha R^{2}$ as an example. As mentioned in the previous section, all coefficients $a_{ij}$, except for $a_{10}=1$ and $a_{2N}=\alpha$, vanish. The first term in the Galilean limit has $i=1$, $j=0$ and $l=2$ i.e, $n=0$. Solving $2n+m+\tfrac{l-j}{2}=i$ leads to $m=0$. The second term has $i=2$, $j=N$, $l=N+4$ i.e, $n=0$. Solving $2n+m+\tfrac{l-j}{2}=i$ gives $m=0$. Putting all together, we get
@@
For this theory to be a modification of the Galilean limit of GR at some order, we must impose the condition $N\leq-2$. Otherwise it would be a theory of pure quadratic gravity.
@@
with the condition $N\leq-2$ to get a GR modification. From this condition and the condition for the Carrollian case, we see that no $f(R)$ theory with a polynomial function $f$ can be a modification of GR in both the Galilean and the Carrollian regimes at the same time.
@@
The NLO, however, is a little different, but the general formula can be derived in a similar way to the Carrollian case,
@@
Let us again check these expressions by applying to GR again. The LO reduces to
@@
The $n-$th order in the Galilean expansion can be derived from the same maximization problem by replacing $l$ with $l-2n$.
@@ IV.2 $f(g_{\mu\nu},R_{\mu\nu\sigma\rho})$ gravity
We first consider gravity theories of the form $f(g_{\mu\nu},R_{\mu\nu\sigma\rho})$ where $f$ is polynomial. We consider a generic term in the Lagrangian to apply the algorithm, the $n$-th term in the Carrollian or Galilean expansions is the sum of the contributions of all terms. A generic term in the Lagrangian can be written as
@@
where $\mathcal{R}$ can be the Riemann tensor, the Ricci tensor with all indices down, or the Ricci scalar, $\bar{\mathcal{R}}$ can be the same tensors with all indices up, and $a$, $b$ are positive numbers such that the Lagrangian terms are scalars. Using eq.(2.8) in [85] and raising/lowering the indices, we can write the generic expansions
$\mathcal{R}$ $=\tfrac{1}{c^{2}}\mathcal{R}_{1}+\mathcal{R}_{2}+c^{2}\mathcal{R}_{3}+c^{4}\mathcal{R}_{4},$$\hskip 10.0pt\bar{\mathcal{R}}$ $=\tfrac{1}{c^{4}}\bar{\mathcal{R}}_{1}+\tfrac{1}{c^{2}}\bar{\mathcal{R}}_{2}+\bar{\mathcal{R}}_{3}+c^{2}\bar{\mathcal{R}}_{4}.$ (67)
The explicit expressions are in the appendix. Raising the expansions to the powers of $a$, $b$, and $c$, respectively, we obtain
@@
Assuming $\alpha=c^{N}\alpha^{\prime}$ where $\alpha^{\prime}$ has no $c$ dependency and substituting in the Lagrangian, we find
@@
The $n$-th order in the Galilean or the Carrollian limits can be obtained by solving a numerical optimization problem as we will demonstrate below.
@@ IV.2.1 Galilean limit
This is a constrained numerical optimization problem that can be solved by the simplex method [88] for $k_{i}$, $\bar{k}_{j}$ given that $a$, $b$ are known. The solution is $z=4a+4b+2+N$. Thus, for this term to be a modification of GR, the condition on $N$ is $N\leq-4a-4b$. If such term exists in the Lagrangian then the Lagrangian is a modification of the Galilean limit of GR. The $n$-th order can be deduced by modifying the constrained optimization problem by $z\to z-2n$.
@@ IV.2.2 Carrollian limit
$\sum_{i=0}^{4}k_{i}=a,$ $\hskip 10.0pt\sum_{j=0}^{4}\bar{k}_{j}=b,$ $k_{i},\bar{k}_{j}\geq 0,$ (73)
which is again solvable by the simplex method given that $a,b$ are known. The solution is $z=N-2a-4b-2$. Thus, for the term to be a modification of GR, the condition on $N$ must be $N\geq 2a+4b$. If such term exists in the Lagrangian then the Lagrangian is a modification of the Carrollian limit of GR. The $n$-th order is given by the modified optimization problem with $z\to z+2n$.
We can see that the conditions on $N$ in the two limits are mutually exclusive. Thus, no such term (of the form (66)) with finite $a$, $b$ can be a GR modification in the two limits simultaneously. This implies that no Lagrangian of the form $f(g_{\mu\nu},R_{\mu\nu\sigma\rho})$ with polynomial $f$ can be a modification of GR in both limits simultaneously. In both limits we can recover the results from the previous section if we set =¯$\mathcal{R}=\bar{\mathcal{R}}=R$. We can recover quadratic gravity examples in the previous section by setting $a=2,b=0$ or $a=0,b=2$ or $a=b=1$.
@@ IV.3 $f(g_{\mu\nu},R_{\mu\nu\sigma\rho},\nabla_{\mu})$ gravity
Now we consider the most general HDG theory, i.e., $f(g_{\mu\nu},R_{\mu\nu\sigma\rho},\nabla_{\mu})$, where $f$ is polynomial. Following [89], any such theory can be recast to the form
@@
where $\mathcal{P}_{0}$ and $\mathcal{Q}_{iI}$ are tensors made up of the Riemann tensor and the metric. The symbols $O_{iI}$ denote differential operators that are constructed from contractions of covariant derivatives. Notice that $\mathcal{P}_{0}$ is just $f(g_{\mu\nu},R_{\mu\nu\sigma\rho})$ which was discussed in the previous subsection. The second term of (74) can be rearranged and written in our notation as
@@
where we added a constant with no $c$ dependency $\alpha^{\prime}$. Moreover, here $i_{\textrm{max}}$, $j_{\textrm{max}}$, $a_{i}$, $b_{ij}$, $c_{ij}$ and $N$ are integers. Expanding the terms using (67) we get
@@
Putting everything together, we arrive at the action
@@
As before, we will now discuss how the Galilean and Carrollian expansions can be transformed into constrained optimization problems.
@@ IV.3.2 Carrollian limit
If such terms exist, then the Lagrangian is a modification of GR in both limits simultaneously. We have shown that theories of the form (66) as well as quadratic gravity do not satisfy these conditions. We will leave the search for such theory (if any exists) to future work.
V Conclusions
In this paper, we reviewed the methods used to construct non-Lorentzian gravitational theories from Lorentzian ones. We showed that all methods of taking the non-Lorentzian limits lead to the same metric, thus, the same non-Lorentzian theories. In the case of the Galilean expansion, the GAG procedure gives the same Galilean theories as the PNR approach. However, the PNR approach explore more theories from expanding the PNR quantities in powers of $c^{-2}$. Such expansion gives rise to infinite number of theories, one for each order, which are not accounted for in other methods. However, these theories are not Galilean till the NLO. The same Galilean theories can be deduced by performing the ADM decomposition and taking the infinite limit of the norm of the orthogonal vector, the IS approach. A similar situation occurs in the Carrollain expansion, although the ZS approach where the limit of the norm of the orthogonal vector is sent to zero (which give the same Carrollian theories as the Carroll group gauging approach) is computationally easier than the PUL approach. The PUL approach explores a larger space of theories, the Carrollian theories to the NLO coincide with the theories we get from other approaches by truncating the expansions of the PUL quantities setting all higher order fields to zero.
Having established that all approaches give the same metric, one can write general expansions for the curvature tensors. We explored the most general HDG theory with polynomial $f$, introducing an algorithm to calculate the $n$-th order of its Galilean and Carrollian expansions as follows:
@@ V Conclusions
    We write the Lagrangian and identify which form it takes (is it of the form (42), (66), or the general (77)) and if the Lagrangian depends on $c$ explicitly.
@@
Another interesting future direction of research is to search for gravity theories which are viable modifications of GR in both Carrollian, and Galilean limits at the same time i.e, satisfies the conditions (82). If such a theory exists, it is interesting to see what significance it has. Is there a defining property of the theory that allows this? If so, how does it impact its solutions. It is also interesting to apply the algorithm introduced to study its Galilean and Carrollian limits, and see how the GR black holes get modified.